# Diversity as a Reward: Fine-Tuning LLMs on a Mixture of Domain-Undetermined Data

### v1

@@ 5.2 Training for Self-Rewarding Abilities
Data Selection: After training the module $\psi_{\text{div}}$, we can use its output to select data samples. Building on the theoretical insights in Sec. 4.4, data points that are closer to other centroids and more dispersed within their own centroid are more beneficial for enhancing the comprehensive capabilities of the model. Therefore, we use the predicted entropy-diversity score as a reward, selecting the top 20% with the highest scores as the final data subset for fine-tuning.
5.3 Empirical Validation and Main Results
To validate the efficacy of DaaR, we conduct experiments comparing more SOTA methods on the data pools in Sec. 3.1 with critical modifications: all domain-specific labels are deliberately stripped. This constraint mimics more challenging real-world scenarios and precludes direct comparison with data mixture methods requiring domain label prior.
Baselines
We use the following data selection methods for comprehensive evaluation: (1) Random Selection: traditional random sampling; (2) Instruction Len: measuring instruction complexity by token count  (Cao et al., 2023); (3) Alpagasus (Chen et al., 2024d): using ChatGPT for direct quality scoring of instruction pairs; (4-5) Instag (Lu et al., 2024): semantic analysis approach with Instag-C (complexity scoring via tag quantity) and Instag-D (diversity measurement through tag set expansion); (6) SuperFilter (Li et al., 2024a): response-loss-based complexity estimation using compact models; (7-9) Deita (Liu et al., 2024b): model-driven evaluation with Deita-C (complexity scoring), Deita-Q (quality scoring), and Deita-D (diversity-aware selection). Detailed configurations for these baselines are in Appendix B.4.
5.3.1 Overall Performance
@@
The experimental results are presented in Table 2, where Instag-Best and Deita-Best represent the optimal variants from their respective method families. Our experiments clearly demonstrate the effectiveness of DaaR across three major language models and seven challenging benchmarks. A detailed analysis is provided below.
@@
High-Difficulty Scenario: The task of balanced capability enhancement proves particularly challenging for existing methods. While some baselines achieve strong performance on specific tasks (e.g., SuperFilter’s 40.55 on HumanEval for Llama3.1), they suffer from catastrophic performance drops in other domains (e.g., SuperFilter’s 6.05 on MATH). Only three baseline methods perform better than Random selection, with notably severe degradation applied in Qwen-series models, all of which fall below the Random performance. We hypothesize this stems from over-specialization – excessive focus on narrow capability peaks at the expense of broad competence (visualized in Appendix B.5). Our method’s robustness stems from preventing extreme distribution shifts through diversity constraints.
@@
Universal Superiority: DaaR establishes new SOTA averages across all models, surpassing the best baselines by +0.14 (Llama3.1), +0.97 (Qwen2), and +1.11 (Qwen2.5). The proposed method uniquely achieves dual optimization in critical capabilities: Mathematical Reasoning: Scores 38.1 MATH (Qwen2) and 16.70 MATH (Qwen2.5), with 7.4% and 27.0% higher than respective random baselines. Coding Proficiency: Maintains 64.94 HumanEval (Qwen2) and 64.20 MBPP (Qwen2.5) accuracy with <1% degradation from peak performance. This demonstrates DaaR’s ability to enhance challenging STEM capabilities while preserving core competencies, a critical advancement over specialized but unstable baselines.
Cost-Efficiency and Flexibility: Compared to baseline methods requiring GPT-based evaluators (Alpagasus, Instag) or full LLaMA-7B inference (Deita), DaaR achieves superior efficiency through data-model co-optimizations. Our method makes the LLM ability of self-rewarding from dedicated data synthesis, and operates on frozen embeddings from layer 3 (vs. full 32-layer inference in comparable methods), reducing computational overhead while maintaining capability integrity. The lightweight 5-layer MLP probe module requires only 9GB GPU memory during training (vs. 18$\sim$24GB for LLM-based evaluators) and adds merely 76M parameters. This plug-and-play feature enables seamless integration of DaaR with existing LLMs without additional dependency management.
6 Conclusion and Future Work
@@
In this paper, we propose a new approach to fine-tuning LLMs with data that lacks clear domain labels, using diversity as a guiding principle. Our method allows LLMs to automatically select and benefit from diverse datasets. By measuring semantic diversity with entropy, we employ a self-reward mechanism built upon the given LLM, identifying data that best fits the model’s natural tendencies in terms of its underlying knowledge distribution. Our experiments with various SOTA LLMs show notable superiority of the method over SOTA methods, highlighting the potential of data diversity to enhance model overall performance.
This research demonstrates feasibility of reward modeling on LLM data diversity and deepens the understanding of its utilities, especially when domain labels are missing or uncertain. Future work could include developing methods to efficiently adjust diversity measures and learning algorithm, thus supporting LLMs in self-evolving towards artificial general intelligence in dynamic environments.
@@
  • Sec. B: The details about experiments and DaaR implementation, including the construction of data pools (Sec. B.1), the benchmarks (Sec. B.2), the training & evaluation implementation (Sec. B.3), the implementation of baselines (Sec. B.4), the visualization of baseline’s selected data samples (Sec. B.5), the implementation of the synthetic data generation (Sec. B.6), the implementation of seed generation & clustering (Sec. B.7), and the prompts for the generation stage (Sec. B.8).
  • Sec C: The full proof for theoretical results are presented in Section 4.
  • Sec D: The more experimental results, including the visualization results of embeddings on Llama3.1-8B & Qwen2.5-7B (Sec D.1), the visualization results of inter- & intra-diversity distributions (Sec D.2), the similarity analysis in generated centroid quantities (Sec. D.3), the similarity analysis in generated centroid domains (Sec. D.4), the DaaR layer selection protocol (Sec. D.5), the comprehensive training dynamics (Sec. D.6), the validation of inter-diversity and intra-diversity (Sec. D.7) and the complete results of DaaR with baselines on all LLMs (Sec. D.8).
@@ Appendix A Notation
For ease of reading and reference, we present all mathematical symbols used in this paper in Table 3.
@@
$\mathcal{K}$ Set of labeled domains $h_{k}$ Independent model of domain $k$
$l$ Loss function $\mathcal{L}_{\mathcal{D}_{k}}$ Empirical risk
$\tilde{\mathcal{D}}_{m}$ Latent distribution of foundational capabilities $\pi_{km}$ Mixture weights
$h_{\theta_{m}}$ Component model $\theta_{m}^{*}$ Optimal predictor
@@
$\mathcal{M}_{\text{ebd}}$ Embedding layer of LLM $\mathcal{M}_{3}$ Layer-3 of LLM
@@
$\psi_{\text{dom}}$ A MLP-based domain classifier $\psi_{\text{div}}$ A MLP-based entropy predictor
@@
Appendix B Experiments and Method Details
@@ B.1 Construction of Data Pool
  • Common Sense: Dolly-15K (Conover et al., 2023) with 15,011 samples, an open source dataset of instruction-following records generated by thousands of Databricks employees in several of the behavioral categories.
  • Reasoning: Cot-en (Chung et al., 2024) with 74,771 samples, is created by formatting and combining nine CoT datasets released by FLAN.
  • Mathematics: Math-Instruct (Yue et al., 2024) with 262,039 samples, is compiled from 13 math rationale datasets, six of which are newly curated by this work. It uniquely ensures extensive coverage of diverse mathematical fields.
  • Coding: Code-Alpaca (Chaudhary, 2023) with 20,016 samples, is constructed from real-world code examples, providing a rich set of tasks designed to guide models in generating accurate and functional code.
@@
B.2 Benchmarks
To evaluate the models’ true capabilities and performance across different domains, we follow the approach of major open-source LLMs (e.g., the Llama3 series (Dubey et al., 2024) and Qwen2 series (Yang et al., 2024)) and select the following widely used evaluation sets. All evaluations were conducted on the OpenCompass platform [^2] https://opencompass.org.cn/.
  • Common Sense: NQ (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017), which cover factual knowledge-based questions of varying difficulty.
  • Reasoning: HellaSwag (Zellers et al., 2019), which effectively evaluates the model’s comprehensive reasoning ability.
  • Mathematics: GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks, which encompass problems ranging from elementary to competition-level difficulty.
  • Coding: MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021), which include evaluations of basic coding abilities in Python. We use the average of various metrics to demonstrate the models’ overall performance across different domains.
@@ B.2 Benchmarks
B.3 Training and Evaluation Details
@@
We implement our approaches using PyTorch (Paszke et al., 2019) v2.4.1, coupled with PEFT v0.12.0 and the Transformers library (Wolf et al., 2020) v4.45.2. Experiments are conducted on a computing platform equipped with four NVIDIA A100 GPUs (40GB), with pre-trained LLMs loaded as 16-bit floating-point numbers. The specific data-model development processes are completed in Data-Juicer Sandbox (Chen et al., 2024c), via integration with the ms-swift (Zhao et al., 2024b) training repository, and the OpenCompass (Contributors, 2023) evaluation repository.
@@
In our experimental setup, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021) adapters for the fine-tuning process, utilizing a LoRA-rank of 8 and a LoRA-alpha of 16. The learning rate was consistently maintained at $5\times 10^{-5}$ across all experiments to ensure uniformity in training dynamics. We utilize a batch size of 4 and set the maximum sequence length to 2048 tokens to accommodate the model’s capacity. To optimize the training process, a warmup ratio of 0.05 was applied and a validation ratio of 0.03 was used. The training was conducted over a single epoch, balancing computational efficiency with the need for effective model adaptation. Following some effective instruction-tuning work (Zhou et al., 2024; Lu et al., 2024), we set the size of our subset to 8,000 entries, which constitutes 20% of the data pool.
@@
Following the guidelines provided by OpenCompass (Contributors, 2023), we adhered to the default settings for our evaluation process. We select the hf-type as base and utilized a batch size of 16 to ensure efficient processing. For most tasks, we employ the gen mode, while for the Hellaswag task, we opt for the ppl mode to better assess perplexity.
B.4 Implementation of Baselines
Random Selection(Rand)
Instruction Length (IL)
The length of the instruction can be considered a measure of input complexity. It is widely believed (Cao et al., 2023; Zhao et al., 2023) that more complex data is beneficial for enhancing model capabilities. Therefore, we select a subset of 8,000 entries with the maximum number of words (based on spaces) in the concatenation of Instruction and Input as part of the filtering process.
AlpaGasus (Chen et al., 2024d)
[…]
Instruction: [Instruction]
Input: [Input]
Response: [Response]
User Prompt:
Figure 3: The distribution of the score by Alpagasus.
Instag (Lu et al., 2024)
[…]
[begin]
{Instruction + Input}
[end]
User Prompt:
Please provide coarse-grained tags, such as ”Spelling and Grammar Check” and ”Cosplay”, to identify main intentions of above instruction. Your answer should be a list including titles of tags and a brief explanation of each tag. Your response have to strictly follow this JSON format: [”tag”: str, ”explanation”: str]. Please response in English.
Figure 4: The distribution of the tags by Instag.
SuperFilter (Li et al., 2024a)
Deita (Liu et al., 2024b)
B.5 Visualization of Baseline’s Selected Data
We project all baseline-selected data samples through Llama3.1-8B’s, Qwen2-7B’s and Qwen2.5-7B’s embedding layer and visualize their distributions via t-SNE dimensionality reduction in Fig. 5, Fig. 6 and Fig. 7 respectively. This visualization provides intuitive insights into how different data selection strategies handle domain-mixed data. Three key observations emerge:
  • 1
    Certain methods (notably Instruction Length, Deita-C, and Deita-Q baselines) exhibit significant distributional bias against coding-domain samples, correlating with their poor HumanEval performance in Table 10.
  • 2
    The visualization reveals inherent challenges in maintaining original data distributions when handling domain-undetermined samples.
  • 3
    Despite observable distribution shifts, not all methods show proportional performance degradation, aligning with our theoretical analysis in Section 4 about the link between latent distribution and LLM’s fundamental capabilities.
Figure 5: t-SNE visualization of data samples selected by different baselines using Llama3.1-8B embeddings. While original labels were removed during training, we preserve them in the legend for interpretability.
Figure 6: t-SNE visualization of data samples selected by different baselines using Qwen2-7B embeddings. While original labels were removed during training, we preserve them in the legend for interpretability.
Figure 7: t-SNE visualization of data samples selected by different baselines using Qwen2.5-7B embeddings. While original labels were removed during training, we preserve them in the legend for interpretability.
B.6 Implementation of the Synthetic Data Generation
@@
B.7 Implementation of Seed Generation & Clustering
Due to the difficulty of pre-trained LLMs in providing accurate and coherent responses, we utilize their corresponding Instruct versions for centroid data generation. Specifically, we apply Llama3.1-8B-Instruct for Llama3.1-8B, Qwen2-7B-Instruct for Qwen2-7B, and Qwen2.5-7B-Instruct for Qwen2.5-7B. The domain description and the zero-shot prompt used in the Seed Generation phase are as follows. When clustering reward training data based on the seed, we select 5,000 entries that match the distribution of the fine-tuning data but do not overlap with it. This practice has proven that even a small amount of data can effectively train a reward probe module of DaaR.

### v2

@@ 5 Experiments
5.1 Overall Performance
@@
The main experimental results are presented in Table 2, where Instag-Best and Deita-Best represent the optimal variants from their method families. Our experiments clearly demonstrate the effectiveness of DaaR across three major language models and seven challenging benchmarks:
High-Difficulty Scenario
@@
: The task of balanced capability enhancement proves particularly challenging for existing methods. While some baselines achieve strong performance on specific tasks (e.g., SuperFilter’s 40.55 on HumanEval for Llama3.1), they suffer from catastrophic performance drops in other domains (e.g., SuperFilter’s 6.05 on MATH). Only 3 baselines perform better than Random on Llama3.1-8B. In the Qwen series, even all baselines fall below the Random performance. We conjecture this stems from selected distribution-bias, as the baselines are unable to balance the proportions across different domains, a perspective visually supported and analyzed in Appendix D.2.
Comprehensive Performance
@@
: DaaR establishes new SOTA averages across all models, surpassing the best baselines by +0.14 (Llama3.1), +0.97 (Qwen2), and +1.11 (Qwen2.5). The proposed method uniquely achieves dual optimization in critical capabilities: Mathematical Reasoning: Scores 38.1 MATH (Qwen2) and 16.70 MATH (Qwen2.5), with 7.4% and 27.0% higher than respective random baselines. Coding Proficiency: Maintains 64.94 HumanEval (Qwen2) and 64.20 MBPP (Qwen2.5) accuracy with <1% degradation from peak performance. This demonstrates DaaR’s ability to enhance challenging STEM capabilities while preserving core competencies.
5.2 Generalizability and Uniqueness of DaaR
Evaluation on MMLU In the main experiments, the data and benchmarks we employed are meticulously aligned, which validates the In-Distribution capabilities of DaaR. To further validate the generalization ability of DaaR, we present MMLU mmlu2021 evaluations in Appendix F.1. MMLU encompasses a significantly broader range of content than our dataset, thus serving as an Out-of-Distribution scenario. Results demonstrate that DaaR achieved top accuracy on Qwen2 with an improvement of 0.61 points, and secured a competitive second-place performance on Qwen2.5. At a granular level, DaaR particularly excels in the STEM domain, outperforming baselines by an average of 7.6% and 1.7% on Qwen2 and Qwen2.5, which aligns the findings above.
Evaluation on Qwen3 In Appendix F.2, we explore the performance of DaaR on the mainstream RL-based LLM, Qwen3-8B. We observe that despite difference in structure, the phenomena (Fig. 8 & Table 8) and the dynamics of training (Fig. 9) remain consistent with the main observations. Then we compare DaaR with baselines in Table 9, although Qwen3-8B is not directly designed for SFT paradigm, DaaR maintains the top position in Avg performance across baselines, surpasses all baselines up to 4.7% and notably exceeds the avg of baselines by 25.91% on MATH.
Customized Characteristics Unlike baselines, DaaR enables domain-aware data selection in undetermined scenario, showcasing potential customization on $\lambda_{k}$. As detailed in Appendix F.3, by adjusting the selection ratio for specific domains, we achieve SOTA performance in targeted domain without compromising overall ability, promising practical utility in real-world application.
Table 3: Illustration results on MMLU, Qwen3 and customization, detailed in Table 6, 9, 10
(a) MMLU Performance
Qwen2-7B MMLU-Avg
Raw 28.98
Rand (8K) 68.81
Instruction-L 68.55
Alpagasus 59.34
Instag-Best 68.65
SuperFilter 54.27
Deita-Best 68.08
DaaR (Ours) 69.42
(b) Eval on Qwen3
Qwen3-8B Avg
Raw 49.41
Rand (8K) 48.16
Instruction-L 49.19
Alpagasus 47.81
Instag-Best 48.69
SuperFilter 49.51
Deita-Best 49.90
DaaR (Ours) 50.06
(c) Customized Selection on Domain
Qwen2.5 NQ Hellaswag Avg
Raw 8.84 72.75 44.64
Rand 11.46 73.08 52.66
Instag 11.08 72.79 52.63
DaaR 15.83 72.48 53.76
Common 17.56 72.51 52.85
Reason 16.56 74.62 53.39
5.3 Ablation Studies and Discussions
Distinct Domain Centroid As shown in Appendix G.1, DaaR can generate domain-representative samples with clear distinctions. Notably, the diversity is consistently most pronounced between common sense, reasoning, and coding domains across model architectures and parameters.
Impacts of Hyper-Parameters In Appendix G.2-Appendix G.6, we conduct a series of ablation studies, including (1) selection of layer: the discrimination accuracy derived from layer 1-10 fluctuates in 91.2-93.6%; (2) the number of seed: the similarity between centroids remains highly consistent with variance ranging from 5.4E-7 to 1.8E-5; (3) the number of augmented data: the mean similarity of centroids generated from sizes of 10 and 30 is 0.983 and 0.981 on Qwen2.5 and Llama3.1 respectively; (4) the size of sliding window & similarity threshold: the impact of different hyper-parameters on centroid generation is minimal, ensuring the subsequent training stability of DaaR.
Cost-Efficiency and Flexibility As shown in Appendix G.7, our method demonstrates significant computational efficiency compared to baselines that rely on GPT-based evaluators or full-LLM inference. By operating on frozen hidden layers and generating regression scores instead of full vocabulary distributions, our approach achieves 70% lower GPU usage and 2.5x faster inference.
6 Conclusions, Limitations & Future Works
@@
In this paper, we propose a new approach to fine-tuning LLMs with data that lacks clear domain labels, using diversity as a guiding principle. By measuring semantic diversity with entropy, we employ a self-reward mechanism built upon the given LLM, identifying data that best fits the model’s natural tendencies in terms of its underlying knowledge distribution. Our experiments with various SOTA LLMs show notable superiority of the method over competitive baselines, highlighting the potential of data diversity to enhance model overall performance.
There are several limitations to our study. Regarding the implementation of DaaR, the customization of selection ratios holds potential for broader application in real-world scenarios. Future work could include developing methods to efficiently adjust diversity measures and learning algorithms, thus supporting LLMs in self-evolving towards artificial general intelligence in dynamic environments.
@@
  • Sec. B: The details about experimental setup, including data pool construction (Sec. B.1), implementation of embedding (Sec. B.2), benchmarks (Sec. B.3) and train / eval details (Sec. B.4).
  • Sec. C: The implementation details of DaaR, including hyper-parameter setting (Sec. C.1) and the process of model-aware training data generation (Sec. C.2).
  • Sec. D: The details of utilized baselines, including the implementation of baselines (Sec. D.1) and visualization of the data selected by baselines (Sec. D.2).
  • Sec. E: The detailed derivation of Proposition 3.2.
  • Sec. F: More comprehensive experiments about DaaR, including the evaluation on MMLU (Sec. F.1), the validation on Qwen3 (Sec. F.2), the customized selection (Sec. F.3) and the utilization of entropy (Sec. F.4).
  • Sec. G: Complete ablation studies and discussions of DaaR, including the centroid similarity (Sec. G.1), LLM’s layer selection (Sec. G.2). As well as several ablation studies of the number of seeds (Sec. G.3), the number of diversity augmentation data (Sec. G.4), the size of sliding window (Sec. G.5) and diversity threshold (Sec. G.6). Additionally, we verify the cost-efficiency (Sec. G.7) and the stability of DaaR (Sec. G.8).
  • Sec H: The comprehensive experimental results, including the visualization results of embeddings on Qwen2-7B, Llama3.1-8B & Qwen2.5-7B (Sec H.1), the visualization of inter- & intra-diversity distributions (Sec H.2), the comprehensive training dynamics (Sec. H.3), the validation of inter-diversity and intra-diversity (Sec. H.4) and complete results of DaaR with baselines (Sec. H.5).
@@ Appendix A Notation
For ease of reading and reference, we present the mathematical symbols used in this paper in Table 4.
@@
$l$ Loss function $p,q$ Data distributions
@@
$\mathcal{M}_{\text{ebd}}$ Embedding layer of LLM $\mathcal{M}$ Hidden layer of LLM
@@
$\psi_{\text{dom}}$ A MLP-based domain predictor $\psi_{\text{div}}$ A MLP-based entropy predictor
@@
Appendix B Experimental Setup Details
@@ B.1 Construction of Data Pool
  • Common Sense: Dolly-15K [14] with 15,011 samples, an open source dataset of instruction-following records generated by thousands of Databricks employees in several of the behavioral categories.
  • Reasoning: Cot-en [12] with 74,771 samples, is created by formatting and combining nine CoT datasets released by FLAN.
  • Mathematics: Math-Instruct [49] with 262,039 samples, is compiled from 13 math rationale datasets, six of which are newly curated by this work. It uniquely ensures extensive coverage of diverse mathematical fields.
  • Coding: Code-Alpaca [6] with 20,016 samples, is constructed from real-world code examples, providing a rich set of tasks designed to guide models in generating accurate and functional code.
@@
B.2 Implementation of Embedding
Given the Alpaca-based [38] data format employed in our study, we concatenate the ’instruction’, ’input’, and ’output’ components of each data sample to capture domain-aware semantic representations. These concatenated sequences are then fed into the LLM, where we compute the average of all token embeddings extracted from the embedding layer to derive the semantic vector characterizing each sample.
The embedding layer is deliberately selected due to its shallow architectural position: (1) it preserves precise semantic information capture through raw input representations, and (2) it incurs reduced computational overhead compared to deeper layers. As evidenced by the visualization in Fig. 12 (Appendix G.2), this approach demonstrates robust domain discrimination capabilities across different LLM layers, thereby validating the robustness of our methodology.
B.3 Benchmarks
To evaluate the models’ true capabilities and performance across different domains, we follow the approach of major open-source LLMs (e.g., the Llama3 series [17] and Qwen2 series [46]) and select the following widely used evaluation sets. All evaluations were conducted on the OpenCompass platform [^1] https://opencompass.org.cn/.
  • Common Sense: NQ [25] and TriviaQA [24], which cover factual knowledge-based questions of varying difficulty.
  • Reasoning: HellaSwag [50], which effectively evaluates the model’s comprehensive reasoning ability.
  • Mathematics: GSM8K [13] and MATH [22] benchmarks, which encompass problems ranging from elementary to competition-level difficulty.
  • Coding: MBPP [4] and HumanEval [11], which include evaluations of basic coding abilities in Python. We use the average of various metrics to demonstrate the models’ overall performance across different domains.
@@ B.3 Benchmarks
B.4 Training and Evaluation Details
@@
We implement our approaches using PyTorch [31] v2.4.1, coupled with PEFT v0.12.0 and the Transformers library [42] v4.45.2. Experiments are conducted on a computing platform equipped with four NVIDIA A100 GPUs (40GB), with pre-trained LLMs loaded as 16-bit floating-point numbers. The specific data-model development processes are completed in Data-Juicer Sandbox [9], via integration with the ms-swift [54] training repository, and the OpenCompass [15] evaluation repository.
@@
In our experimental setup, we employ Low-Rank Adaptation (LoRA) [23] adapters for the fine-tuning process, utilizing a LoRA-rank of 8 and a LoRA-alpha of 16. The learning rate was consistently maintained at $5\times 10^{-5}$ across all experiments to ensure uniformity in training dynamics. We utilize a batch size of 4 and set the maximum sequence length to 2048 tokens to accommodate the model’s capacity. To optimize the training process, a warmup ratio of 0.05 was applied, and a validation ratio of 0.03 was used. The training was conducted over a single epoch, balancing computational efficiency with the need for effective model adaptation. Following some effective instruction-tuning work [55, 30], we set the size of our subset to 8,000 entries, which constitutes 20% of the data pool.
@@
Following the guidelines provided by OpenCompass [15], we adhered to the default settings for our evaluation process. We select the hf-type as base and utilize a batch size of 16 to ensure efficient processing. For most tasks, we employ the gen mode, while for the Hellaswag task, we opt for the ppl mode to better assess perplexity.
Appendix C Implementation of DaaR
C.1 Implementation of the Synthetic Data Generation
@@
For implementation details of DaaR’s hyper-parameters, we set the number of seed samples $\mathcal{S}_{k}^{(0)}$ as 5 (ablation in Appendix G.3), the number of anchors in sliding window as 3 (ablation in Appendix G.5), the similarity threshold $\tau$ as 0.9 for Mathematics and 0.85 for others (ablation in Appendix G.6), the terminated iteration as 30 (ablation in Appendix G.4), the hidden layer $\mathcal{M}(\tilde{x})$ as layer-3 (ablation in Appendix G.2).
C.2 Implementation of Model-Aware Training Data Generation
Seed Generation
This phase aims to provide zero-shot LLM-generated seed data that reflects how LLMs inherently perceive the characteristics of domain-specific data. We employ the LLM itself to generate concise domain descriptions across diverse fields illustrated in Prompt C.2, and the prompt shown in Prompt C.2. Due to the difficulty of pre-trained LLMs in providing accurate and coherent responses, we utilize their corresponding Instruct versions for centroid data generation. Specifically, we apply Llama3.1-8B-Instruct for Llama3.1-8B, Qwen2-7B-Instruct for Qwen2-7B, and Qwen2.5-7B-Instruct for Qwen2.5-7B.
Notably, these descriptions are not required to be optimal or prescriptive, rather, they serve as representative examples compatible with our experimental scenarios. Users retain the flexibility to customize target domain specifications according to specific requirements. The associated stability analysis regarding this design choice will be elaborated in Appendix G.8.
Diversity Augmentation
While seed data effectively captures the intrinsic domain knowledge of LLMs, its limited quantity and the constrained diversity of single-prompt generation methods necessitate enhancement. We address these limitations through two complementary strategies including a sliding window mechanism and a rule-based filter. The implementation details are formalized in Prompt C.2 and Eq. (7).
Domain-Aware Clustering
During reward model training data clustering, we rigorously curate a subset of 5,000 entries that are distribution-matched to the fine-tuning dataset while separate from it.